0 0
Add a comment Improve this question Transcribed image text
Answer #1

xcm =17.6211cm

ycm =18.1067cm
Answer #2

Size of 1 block of chessboard= 33.3/8 cm= 4.1625cm

x-coordinate=

85.5*4*4.1625+18.7*(1.5+3.5+5.5+7.5)*4.1625/(85.5+4*18.7)

=17.621cm

y- coordinate

85.5*4*4.1625+18.7*(2.5+3.5+5.5+7.5)*4.1625/(85.5+4*18.7)

= 18.1067cm